# SU(4) Kondo Lattice in Semiconductor Moiré Materials

### v1

@@ Abstract
Given the manifestations of rich Kondo physics in TMD moiré materials, it is natural to explore related platforms that may host even richer phenomena. One promising direction is the realization of Kondo lattices with an enlarged number of degrees of freedom, that is, Kondo lattices exhibiting SU($N>2$) symmetry [29, 30, 31, 32, 33]. Such models are particularly appealing since SU($N>2$) spins feature enhanced spin fluctuations that can stabilize exotic ground states [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. In graphene-based moiré materials, SU(4) and SU(8) Kondo lattice models have been proposed in twisted trilayer [49] and bilayer graphene [50], with the SU(8) case in magic-angle twisted bilayer graphene rooted in a topological heavy fermion picture [51]. In TMD moiré materials, a natural strategy is to exploit layer degrees of freedom [52, 53, 54, 55]. Recently, Coulomb-coupled TMD bilayers with suppressed interlayer tunneling was theoretically proposed as a platform for simulating an SU(4) Hubbard model [53], followed by experimental evidence in twisted AB-stacked homobilayers [56].
In this work, we extend these studies by proposing TMD moiré multilayers as a platform for realizing an SU(4)-symmetric triangular Kondo lattice. Our model goes beyond the conventional Kondo lattice model by incorporating a novel three-site exchange of SU(4) local moments, which induces spontaneous breaking of time-reversal and lattice symmetries. Using a parton mean-field approach, we map out the quantum phase diagram as a function of three-site exchange and doping, as illustrated schematically in Fig. 1(a). Below critical doping, we identify various Mott insulators including a chiral spin liquid and crystallized bond-ordered phases. Above critical doping, a heavy Fermi liquid with broken lattice symmetry emerges, with or without anomalous Hall responses. We expect our mean-field results to be reasonably reliable, since fluctuations around the saddle point are relatively suppressed in the SU(4) case. Our theoretical study highlights that the TMD moiré SU(4) Kondo lattice can serve as a rich playground for realizing a plethora of correlated and topological phases.
@@
Model Hamiltonian.- The model Hamiltonian for these two settings can be written as
@@
$\tilde{H}$ $=$ $-t_{c}\sum_{\langle ij\rangle,\alpha}\left(c^{\dagger}_{i,\alpha}c_{j,\alpha}+\textnormal{h.c.}\right)-t_{f}\sum_{\langle ij\rangle,\alpha}\left(f^{\dagger}_{i,\alpha}f_{j,\alpha}+\textnormal{h.c.}\right),$ (2)
@@
In Fig. 2, we present a mean-field phase diagram as a function of $J_{3}/J_{2}\>(<0)$ and $\nu_{c}$, with other parameters $(t_{c},J_{2},J_{K})$ being fixed. The phase diagram exhibits a density-tuned Kondo transition at critical doping $\nu_{c}$, which depends on $J_{3}/J_{2}$. In the Kondo-unscreened regime ($\Phi\rightarrow 0$), the doped $c$-holes form a homogeneous Fermi liquid with a small Fermi surface, while the $f$-fermions exhibit various Mott insulating phases: a plaquette-order phase, a chiral spin liquid (CSL), and a decoupled chain (DC). These phases are in good agreement with a previous study of the SU(4) Hubbard model [53]. In the Kondo-screened regime ($\Phi>0$), a heavy Fermi liquid appears across a wide range of $\nu_{c}$, and exhibits various lattice symmetry-breaking patterns. Below, we discuss the nature of each phase in detail, beginning with the Kondo-unscreened phases.
@@
For sufficiently negative $J_{3}$, the two-site coupling drives a strong dimerization, leading to the DC phase. We note that the three-site coupling term ($\propto d_{abc}$) survives in the SU($N>2$) case even in the absence of an external magnetic flux, in contrast to the SU(2) case.
@@
Outlook.- In this work, we proposed TMD multilayer settings that realize an SU(4) Kondo lattice and performed a self-consistent mean-field study of its phase diagram. We have focused on a specific choice of parameters that does not necessarily derive from a microscopic model, since our main goal is to outline possible quantum phases that can be realized in this setting. An interesting future direction is to incorporate realistic microscopic parameters to map out the phase diagram as a function of gate voltages, including the effects of RKKY interactions, and to study the robustness of each phase against SU(4)-breaking couplings [66].
@@
Acknowledgements.- I thank T. Senthil, Allan MacDonald, and Jie Shan for stimulating discussions at the Boulder Summer School 2025. I am grateful to Kin Fai Mak and Wenjin Zhao for collaboration on a related project, and especially to Kin Fai Mak for insightful discussions during my A exam that helped motivate this work. I thank Aaditya Panigrahi, Haoyu Guo, Omri Lesser, and Xuepeng Wang for helpful discussions and comments. I am deeply indebted to Debanjan Chowdhury for his guidance throughout my PhD and invaluable discussions related to this work. I acknowledge the support of the group and access to shared computational resources.
@@ I Schrieffer-Wolff transformation
In this section, we provide details of the Schrieffer-Wolff transformation [59, 60]. We begin by rewriting Eq. 2 as
@@ III CSL and DC for $J_{3}/J_{2}<0$
Note that $J_{2}=1$ and $J_{3}=6t^{3}_{f}/U^{2}<0$ in our model. For sufficiently negative $J_{3}$, the first term leads to strong dimerizations. Therefore, the decoupled chain phase can be favorable.

### v2

@@ Abstract
Given the manifestations of rich Kondo physics in TMD moiré materials, it is natural to explore related platforms that may host even richer phenomena. One promising direction is the realization of Kondo lattices with an enlarged number of degrees of freedom, that is, Kondo lattices exhibiting SU($N>2$) symmetry [29, 30, 31, 32, 33]. Such models are particularly appealing since SU($N>2$) spins feature enhanced spin fluctuations that can stabilize exotic ground states [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. In graphene-based moiré materials, SU(4) and SU(8) Kondo lattice models have been proposed in twisted trilayer [50] and bilayer graphene [51], with the SU(8) case in magic-angle twisted bilayer graphene rooted in a topological heavy fermion picture [52]. In TMD moiré materials, a natural strategy is to exploit layer degrees of freedom [53, 54, 55, 56, 57]. Recently, Coulomb-coupled TMD bilayers with suppressed interlayer tunneling was theoretically proposed as a platform for simulating an SU(4) Hubbard model [54], followed by experimental evidence in twisted AB-stacked homobilayers [58].
In this work, we extend these studies by proposing TMD moiré multilayers as a platform for realizing an SU(4)-symmetric triangular Kondo lattice. Our model goes beyond the conventional Kondo lattice model by incorporating a three-site exchange of SU(4) local moments, which induces spontaneous breaking of time-reversal and lattice symmetries. Using a parton mean-field approach, we map out the quantum phase diagram as a function of three-site exchange and doping, as illustrated schematically in Fig. 1(a). Below critical doping, we identify various Mott insulators including a chiral spin liquid and crystallized bond-ordered phases. Above critical doping, a heavy Fermi liquid with broken lattice symmetry emerges, with or without anomalous Hall responses. We expect our mean-field results to be reasonably reliable, since fluctuations around the saddle point are relatively suppressed in the SU(4) case. Our theoretical study highlights that the TMD moiré SU(4) Kondo lattice can serve as a rich playground for realizing a plethora of correlated and topological phases.
@@
Model Hamiltonian.- Motivated by these considerations, we introduce below a minimal Anderson lattice Hamiltonian for the proposed multilayer settings. Closely related lattice models have been derived for TMD trilayers from continuum moiré band structures combined with self-consistent Hartree-Fock calculations [22]. The Hamiltonian below can be understood as a mirror-symmetric two-copy extension of the trilayer building block, which is what we mean above by a generalization of the double quantum dot model. The model Hamiltonian can be written as
@@
$\tilde{H}$ $=$ $-t_{c}\sum_{\langle ij\rangle,\alpha}\left(c^{\dagger}_{i,\alpha}c_{j,\alpha}+\textnormal{h.c.}\right)-t_{f}\sum_{\langle ij\rangle,\alpha}\left(f^{\dagger}_{i,\alpha}f_{j,\alpha}+\textnormal{h.c.}\right)$ (2)
@@
In Fig. 2, we present a mean-field phase diagram as a function of $J_{3}/J_{2}\>(<0)$ and $\nu_{c}$, with other parameters $(t_{c},J_{2},J_{K})$ being fixed. The phase diagram exhibits a density-tuned Kondo transition at critical doping $\nu_{c}$, which depends on $J_{3}/J_{2}$. In the Kondo-unscreened regime ($\Phi\rightarrow 0$), the doped $c$-holes form a homogeneous Fermi liquid with a small Fermi surface, while the $f$-fermions exhibit various Mott insulating phases: a plaquette-order phase, a chiral spin liquid (CSL), and a strongly anisotropic, quasi-1D (“decoupled-chain”, DC) regime. These phases are in good qualitative agreement with a previous DMRG study of the SU(4) Hubbard model with ring exchange [54], which also reports consistency with large-$N$ mean-field studies. In the Kondo-screened regime ($\Phi>0$), a heavy Fermi liquid appears across a wide range of $\nu_{c}$, and exhibits various lattice symmetry-breaking patterns. Below, we discuss the nature of each phase in detail, beginning with the Kondo-unscreened phases. We note that, as a saddle-point treatment, the phase boundaries may shift beyond mean-field, especially in gapless regimes.
@@
For sufficiently negative $J_{3}$, the two-site coupling drives a strong bond ordering, leading to the DC phase. We note that the three-site coupling term ($\propto d_{abc}$) survives in the SU($N>2$) case even in the absence of an external magnetic flux, in contrast to the SU(2) case.
@@
Outlook.- In this work, we proposed TMD multilayer settings that realize an SU(4) Kondo lattice and performed a self-consistent mean-field study of its phase diagram. We have focused on a specific choice of parameters that does not necessarily derive from a microscopic model, since our main goal is to outline possible quantum phases that can be realized in this setting. An interesting future direction is to incorporate realistic microscopic parameters (e.g., extracted from continuum moiré models and self-consistent Hartree-Fock treatments of TMD multilayers) to map out the phase diagram as a function of gate voltages, including the effects of RKKY interactions, and to study the robustness of each phase against SU(4)-breaking couplings [68]. It will also be important to carefully consider layer hybridizations in order to ensure the emergence of well-defined $f$- and $c$-layers [69]. It would also be interesting to work with fully projected (Gutzwiller-projected) wavefunctions, which can quantitatively renormalize energies and shift mean-field phase boundaries, with potentially stronger effects in gapless regimes such as the heavy Fermi liquid.
@@
Acknowledgments.- I thank T. Senthil, Allan MacDonald, and Jie Shan for stimulating discussions at the Boulder Summer School 2025. I am grateful to Kin Fai Mak and Wenjin Zhao for collaboration on a related project, and especially to Kin Fai Mak for insightful discussions during my A exam that helped motivate this work. I thank Aaditya Panigrahi, Haoyu Guo, Omri Lesser, and Xuepeng Wang for helpful discussions and comments. I am deeply indebted to Debanjan Chowdhury for his guidance throughout my PhD and invaluable discussions related to this work. I acknowledge the support of the group and access to shared computational resources.
@@ I Schrieffer-Wolff transformation
In this section, we provide details of the Schrieffer-Wolff transformation. We begin by rewriting Eq. 2 as
@@ III CSL and DC for $J_{3}/J_{2}<0$
Note that $J_{2}=1$ and $J_{3}=6t^{3}_{f}/U^{2}<0$ in our model. For sufficiently negative $J_{3}$, the first term leads to strong bond ordering. Therefore, the decoupled chain phase can be favorable.